# Entanglement corner dependence in two-dimensional systems: A tensor network perspective

### v2

@@ IV Conclusions and future outlook
We show, based on the bond dimension of the iPEPS and its relation to entanglement, that in the continuum limit a corner-dependent and area-independent term emerges, similar to the one predicted in continuous gapped systems. A similar argument shows that such a behavior may also rise from the lattice-corner-dependence of the entanglement in gauge-invariant systems. Our result ties the entanglement behavior in continuous systems with the geometrical behavior of its discretized version in the general case, as was shown in Ref. [42] for the integrated quantum metric of free fermions.
The technical aspects of our result lead to an important practical insight: the behavior observed in the continuous system could not have emerged from a single fixed lattice orientation $\varphi$. It required averaging over all possible orientations to restore the system’s isotropy. Similarly, this term would not appear if the discretized lattice was not translational invariant, e.g., if it were centered with respect to the system’s center. This emphasizes the necessity of such symmetrization procedures when approximating continuous systems by discretization. It would be interesting to try and obtain a similar result for critical systems using discretization on a fractal, obtaining a similar insight for conformal field theories.
@@
Appendix A Bond dimension corner dependence when $\theta\rightarrow\pi$
In this section, we obtain explicitly the dependence of $n_{\text{legs}},n_{\text{corners}}$ on $\theta$ for $\theta\rightarrow\pi$, and show that the dependence matches Eq. (7) as in QFT.
@@ Appendix A Bond dimension corner dependence when $\theta\rightarrow\pi$
We now integrate over the orientation $\varphi$. Note that the expression above is positive for $\varphi-m\pi/2\in(\theta-\pi/2,\pi/2)$ where $m$ is an integer. For orientations outside of this range, $p_{\text{skip}}$ may be geometrically shown to be 0, as demonstrated in Fig. 6c. The final expression is obtained to be
@@
with a corner dependence identical to the continuous system behavior in Eq. (7). The above analysis applies for all $\theta$s large enough such that they will surely pick up the bonds in the neighboring plaquette, which in a square lattice requires $\theta<2\tan^{-1}(1/2)\approx 0.3\pi$. Therefore, Eq. (12) applies to the range $0.3\pi<\theta<\pi$, and not only $\theta\rightarrow\pi$.
We now focus on $n_{\text{corners}}$. In Fig. 6d, we depict the vertex positions in which a boundary corner $\theta$ of orientation $\varphi$ picks up a corner of the lattice. Here we consider different ranges of orientations $\varphi$. The probability to pick up a corner when $\varphi\in(\theta-\pi/2,\pi/2)$ is again obtained geometrically:
@@
Since $\theta\rightarrow\pi$, we obtain
@@
obtaining the behavior of Eq. (7), similarly to $n_{\text{legs}}$. Equivalent calculations may be done for the remaining values of $\varphi$ to obtain a similar dependence on $\pi-\theta$.
Figure 6: (a) When $\theta\rightarrow\pi$, its effect on $n_{\text{legs}}$ stems from the bond legs closest to the vertex. We observe a corner $\theta$ with orientation $\varphi$. Different positions in the vertex result in different contributions to $n_{\text{legs}}$. The red edges crosses the top and bottom bonds, while the blue edges do not pick up any lattice site in the unit cell, i.e., no lattice site is included in $A$. Thus, the edges do not cross any bond. (b) For given $\theta,\varphi$, the blue area denotes the vertex positions which skip a bond. It may be geometrically shown to be $a^{2}/2(\tan\left(\theta-\varphi)+\tan\varphi\right)$. (c) When $\varphi<\theta-\pi/2$, it is geometrically apparent that at least on of the lattice sites is included in $A$ (see the site colored in red). Thus, $p_{\text{skip}}(\theta,\varphi)=0$ in these orientations. (d) For given $\theta,\varphi$, the unit cell subsets which contribute to $n_{\text{corners}}$ are colored in light blue. Example edges are added in dark blue for each of the cases.

### v4

@@ IV Conclusions and future outlook
We show, based on the bond dimension of the iPEPS and its relation to entanglement, that in the continuum limit a corner-dependent and area-independent term emerges, similar to the one predicted in continuous gapped systems. A similar argument shows that such a behavior may also rise from the lattice-corner-dependence of the entanglement in gauge-invariant systems. Our result ties the entanglement behavior in continuous systems with the geometrical behavior of its discretized version in the general case. It could be useful to compare this viewpoint to the one presented in Ref. [45] for the charge fluctuations in free fermion systems, which is related there to the integrated quantum metric.
The technical aspects of our result lead to an important practical insight: the behavior observed in the continuum could not have emerged from a single fixed lattice orientation $\varphi$. It required averaging over all possible orientations to restore the system’s isotropy. Similarly, this term would not appear if the discretized lattice was not translational invariant, e.g., if it were centered with respect to the system’s center. This emphasizes the necessity of such symmetrization procedures when approximating continuous systems by discretization.
@@
V Acknowledgements
We thank E. Bettelheim, P.M. Tam, and E. Zohar for fruitful discussions. Our work has been supported by the Israel Science Foundation (ISF) and the Directorate for Defense Research and Development (DDR&D) through Grant No. 3427/21, the ISF Grant No. 1113/23, and the US-Israel Binational Science Foundation (BSF) through Grant No. 2020072. N.F. is also supported by the Azrieli Foundation Fellows program and the Milner Foundation.
Appendix: Corner dependence of the bond dimension when $\theta\rightarrow\pi$
Figure 6: (a) When $\theta\rightarrow\pi$, its effect on $n_{\text{legs}}$ stems from the bond legs closest to the vertex. We examine a corner with an opening angle $\theta$ and an orientation angle $\varphi$. Different positions of the vertex result in different contributions to $n_{\text{legs}}$. The red edges crosses the top and bottom bonds, while the blue edges do not pick up any lattice site in the unit cell, i.e., no lattice site is included in $A$. Thus, the edges do not cross any bond. (b) For given $\theta,\varphi$, the blue region denotes the vertex positions which lead to skipping of a bond. Its area may be geometrically shown to be $a^{2}/2(\tan\left(\theta-\varphi)+\tan\varphi\right)$. (c) When $\varphi<\theta-\pi/2$, it is geometrically apparent that at least on of the lattice sites is included in AA (see the site colored in red). Thus, $p_{\text{skip}}(\theta,\varphi)=0$ for these orientations. (d) For given $\theta,\varphi$, the unit cell regions which contribute to $n_{\text{corners}}$ are colored in light blue. Example edges are added in dark blue for each of the cases.
In this Appendix, we obtain explicitly the dependence of $n_{\text{legs}},n_{\text{corners}}$ on $\theta$ for $\theta\rightarrow\pi$, and show that the dependence matches Eq. (7) as in continuum field theory.
@@ V Acknowledgements
We now integrate over the orientation $\varphi$. Note that the expression above is positive for $\varphi-m\pi/2\in(\theta-\pi/2,\pi/2)$ where $m$ is an integer. For orientations outside of this range, $p_{\text{skip}}$ may be geometrically shown to be 0, as illustrated in Fig. 6c. The final expression is thus obtained,
$p_{\text{skip}}(\theta)=\int_{\theta-\pi/2}^{\pi/2}p_{\text{skip}}(\theta,\varphi)d\varphi=\frac{1}{4}\left(1-(\pi-\theta)\cot\theta\right),$ (12)
@@
We now turn to on $n_{\text{corners}}$. In Fig. 6d, we depict the vertex positions for which a boundary corner $\theta$ of orientation $\varphi$ picks up a corner of the lattice. Here we consider different ranges of orientations $\varphi$. Let us start from the probability to pick up a corner when $\varphi\in(\theta-\pi/2,\pi/2)$, which is again obtained geometrically:
@@
Since $\theta\rightarrow\pi$, we find
@@
recovering the behavior of Eq. (7), similarly to $n_{\text{legs}}$. Equivalent calculations may be performed for the remaining values of $\varphi$ to obtain a similar dependence on $\pi-\theta$.